# Graph quandles: Generalized Cayley graphs of racks and right quasigroups

### v3

@@ Abstract.
We solve two open problems of Valeriy Bardakov about Cayley graphs of racks and graph-theoretic realizations of right quasigroups. We also extend Didier Caucal’s classification of labeled Cayley digraphs to right quasigroups and related algebraic structures like quandles.
First, we characterize markings of graphs that realize racks. As an application, we construct rack-theoretic (di)graph invariants from permutation representations of graph automorphism groups. We describe how to compute these invariants with general results for path graphs and cycle graphs.
Second, we show that all right quasigroups are realizable by edgeless graphs and complete (di)graphs. Using Schreier (di)graphs, we also characterize Cayley (di)graphs of right quasigroups $Q$ that realize $Q$. In particular, all racks are realizable by their full Cayley (di)graphs.
Finally, we give a graph-theoretic characterization of labeled Cayley digraphs of right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
@@ 1. Introduction
Right quasigroups vastly generalize racks and quandles, which are nonassociative algebraic structures used to construct invariants in group theory, knot theory, and low-dimensional topology. During a 2020 lecture, Bardakov introduced a way to construct right quasigroups from markings of graphs by graph automorphisms [bardakov]. Interested in finding geometric interpretations of racks and quandles, Bardakov closed the lecture by posing two open questions about which marked graphs realize racks and how they relate to Cayley graphs of racks; see Problems 1.6 and 1.7.
In this paper, we answer Bardakov’s questions with the following results. We also introduce two rack-theoretic (di)graph invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ and characterize labeled Cayley digraphs of racks and quandles. We refer to directed graphs as digraphs and simple undirected graphs as graphs.
@@
Let $\Gamma$ be a (di)graph with vertex set $V$, and let $R:V\to\operatorname{Aut}\Gamma$ be a marking (resp. q-marking) of $V$. Then the right quasigroup $V^{\Gamma}_{R}$ realized by $\Gamma$ is a rack (resp. quandle) if and only if $R$ is a magma homomorphism from $V^{\Gamma}_{R}$ to $\operatorname{Conj}(\operatorname{Aut}\Gamma)$. (See Theorem 5.1.)
@@
All right quasigroups are realizable by edgeless graphs and complete (di)graphs. (See Proposition 3.9.)
Theorem 1.3.
@@
All racks are realizable by their full Cayley (di)graphs. (See Corollary 6.4.)
To our knowledge at the time of writing, this paper is the first to study Bardakov’s constructions. It is also the first to characterize labeled generalized Cayley digraphs of right quasigroups, racks, and quandles, as summarized below; cf. Problem 1.8. This extends Caucal’s work on labeled Cayley digraphs of quasigroups, left quasigroups [caucal], groups, monoids, and semigroups [caucal2].
Theorem 1.5.
Let $\mathcal{Q}$ be the class of all labeled digraphs that are deterministic, source-complete, codeterministic, and target-complete. Then $\mathcal{Q}$ is precisely the class of labeled Cayley digraphs of right quasigroups. (See Theorem 7.12.)
Moreover, there exist graph theoretic-conditions on $\mathcal{Q}$ that restrict it to the subclass of labeled Cayley digraphs of racks, and similarly for quandles, involutory racks, and kei. (See Theorem 7.14 and Corollary 7.15.)
1.1. Motivating discussion
We give an overview of the historical background and open questions motivating this work. See Section 2 for formal definitions of the terms discussed here.
1.1.1. Quandle theory and generalizations
[…]
Generalizations of quandles have also been of interest in many areas of mathematics. In particular, Fenn and Rourke [fenn] introduced racks in 1992 to construct complete invariants of closed, connected 3-manifolds and framed links embedded in them. Racks are also used to study Hopf algebras (e.g., [hopf]), the Yang–Baxter equation (see [yang]), and Legendrian knots in contact three-space (see [ta]), for instance. On the other hand, right quasigroups are a class of magmas or groupoids that vastly generalizes racks, quandles, groups, and quasigroups. Various authors have used right quasigroups to study column Latin squares [Latin], smooth deformations of Lie group structures [diff], and nonassociative generalizations of Hopf algebras [quantum], for instance.
1.1.2. Racks as symmetries
Because of the myriad uses of racks and quandles, novel ways of understanding their structure are desirable. Inspired by a class of quandles called dihedral quandles, which model reflections of cycle graphs, Bardakov sought geometric interpretations of other classes of racks and quandles during a 2020 lecture [bardakov]. To formalize this search, he introduced realizations of right quasigroups by marked graphs, which are graphs $\Gamma$ with an assignment $R$ of each vertex to a graph automorphism. Bardakov closed his lecture with the following open problems.
Problem 1.6.
Under what conditions does a marked graph realize a rack or a quandle?
Problem 1.7.
Theorem 5.1 answers Problem 1.6. Since all racks are right quasigroups, Proposition 3.9 answers a generalized version of the first question in Problem 1.7. Corollary 6.4 answers the second question in Problem 1.7, while Theorems 6.16.2 answer generalized forms of the question.
@@ 1.1.3. Characterization of labeled Cayley digraphs
Caucal answered Problem 1.8 for left quasigroups, quasigroups [caucal], semigroups, and various classes of monoids, including groups [caucal2]. [^2] Since the right-multiplication maps $R_{v}$ of left quasigroups are not necessarily permutations, the answers to Problem 1.8 for left and right quasigroups are distinct. Chishwashwa et al. addressed a similar question for vertex-labeled Cayley digraphs of unital, fixed point-free right quasigroups [groupoid1]. In this paper, we answer Problem 1.8 for various classes of racks and right quasigroups; see Theorems 7.12 and 7.14 and Corollary 7.15.
@@ 1.2. Structure of the paper
In Section 5, we answer Problem 1.6. As an application, we introduce two rack-theoretic (di)graph invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ with general results for path graphs and cycle graphs.
In Section 6, we answer Problem 1.7 and its analogues for right quasigroups and digraphs.
In Section 7, we define labeled Cayley digraphs of magmas in the sense of Caucal, and we answer Problem 1.8 for right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
@@ 1.3. Notation
For all positive integers $n\in\mathbb{Z}^{+}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. Denote the symmetric group of $[n]$ by $S_{n}$ with its elements written in cycle notation, and denote the symmetric group of any other set $X$ by $S_{X}$. We also denote the composition of functions $\varphi:V\to W$ and $\psi:W\to X$ by $\psi\varphi$, and we denote the identity map on a set $V$ by $\operatorname{id}_{V}$. For $n\geq 3$, let $D_{n}$ be the dihedral group of order $2n$. Given a subset $S$ of a group $G$, let $S^{-1}:=\{s^{-1}\mid s\in S\}$.
[…]
$v\triangleright w=R_{w}(v).$
@@ 2.1. Magmas
A magma or groupoid is a pair $(V,R)$, denoted by $V_{R}$, where $V$ is a set and $R$ is a mapping from $V$ to the set of functions from $V$ to VV. [^3] Although the notation $V_{R}$ is nonstandard, we use it because Bardakov formulated his open problems with it [bardakov]. For all $v\in V$, we call the map $R_{v}:=R(v)$ a right-multiplication map or right-translation map.
@@
If we take $G:=\mathbb{Q}$ to be the rational numbers in Example 2.2, then the positive rational numbers $\mathbb{Q}^{+}$ are a submagma of $\mathbb{Q}$ but not a right quasigroup.
@@
With this definition, magmas and right quasigroups form categories. In particular, we can consider automorphism groups $\operatorname{Aut}V_{R}$ of magmas.
@@ 2.2.2. Preliminary results
An alternative characterization of racks does the heavy lifting in solving Problem 1.6; see Theorem 5.1.
@@
$\implies$” Suppose that $V_{R}$ is a rack. Then $R(V)\subseteq S_{V}$, and Equation (2.1) states that
@@
[…]
@@ 3. Graph-theoretic preliminaries
We discuss directed and undirected graphs, marked graphs as constructed by Bardakov [bardakov] and Cayley graphs of magmas as introduced by Caucal [caucal]. We also relate the latter to Schreier graphs of group actions. (Since we only discuss labeled digraphs in Section 7, we defer defining them until then.)
@@ 3.1. Graphs and digraphs
We recall the graph-theoretic constructions of Bardakov in [bardakov]. Like Bardakov, we assume that all undirected graphs are simple, and we do not allow for digraphs to have multiple edges. However, we do allow for digraphs to have loops.
@@
  • A digraph or directed multigraph $\Gamma$ is a pair $(V,E)$ where $V$ is a set and $E$ is a subset of $V\times V$. We say that $V$ is the vertex set of $\Gamma$, and we say that $E$ is the (directed) edge set or arc set of $\Gamma$, respectively.
@@ 3.2. Marked graphs
Let $n\in\mathbb{Z}^{+}$ be a positive integer, and let $\Gamma$ be the star graph $K_{1,n}$ of order $n+1$; cf. Table 1 in Section 5. Then $\operatorname{Aut}\Gamma\cong S_{n}$ acts on $V:=V(\Gamma)$ by permuting the leaves.
@@ 3.2.1. Right quasigroups are realizable
We answer the first question in Problem 1.7. Recall that a digraph $\Gamma=(V,E)$ is called complete if
@@ 3.4.1. Preliminary results
Our solutions to Problem 1.7 can be stated nicely in terms of Schreier (di)graphs; cf. Theorems 6.16.2.
@@ 4. Motivating examples
Figure 1 depicts the full Cayley digraph $\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure 1. Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.1.
@@
Let $S=\{1\}$. Figure 2 depicts the partial Cayley digraph $\Gamma(V_{R},S)$, the full Cayley digraph $\Gamma(V_{R})$, and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure 2. Partial and full Cayley digraphs and full Cayley graph of the quandle from Example 4.2.
@@
Nonisomorphic right quasigroups may share the same Cayley graphs and even the same Cayley digraphs. For example, let $V$ be the set $[3]$. Figure 3 depicts the full Cayley digraph of the right quasigroup $V_{R}$ defined by
@@
Figure 3. Full Cayley digraphs of the two right quasigroups from Example 4.3 and their shared full Cayley graph.
@@
Figure 4 depicts the full Cayley digraph $\Gamma:=\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure 4. Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.4.
@@
Let $S=\{1\}$. Figure 5 depicts the partial Cayley digraph $\Gamma=\Gamma(V_{R},S)$ and its underlying graph $\Gamma_{\mathrm{und}}(V_{R},S)$.
@@
Figure 5. Partial Cayley digraph and underlying Cayley graph of the right quasigroup from Example 4.5.
@@ 5. From marked graphs to racks
In this section, we briefly deduce an answer to Problem 1.6. As an application, we introduce two rack-theoretic invariants of graphs.
5.1. Solution to Problem 1.6
@@
Given a (di)graph $\Gamma$, let $\mu_{\mathrm{rack}}(\Gamma)$ (resp. $\mu_{\mathrm{qnd}}(\Gamma)$) be the number of markings of $\Gamma$ that realize racks (resp. quandles). As an application of Theorem 5.1, we compute these numbers for several graphs and discuss how to compute them in general. This is motivated by Problem 1.6.
@@
We provide an implementation of this exhaustive search algorithm in a GitHub repository [code]. With this implementation, we were able to compute $\mu_{\mathrm{rack}}$ and $\mu_{\mathrm{qnd}}$ for complete graphs $K_{n}$ (equivalently, edgeless graphs), star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$; see Table 1. We also have the following general results for path graphs and cycle graphs.
@@
Table 1. Computations of $(\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}})$ for complete graphs $K_{n}$, star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$.
@@
When $n=3$, five markings of the cycle graph $C_{3}$ realize quandles. In particular, let $V_{R}$ be the quandle from Example 4.2. The full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)\cong C_{3}$ depicted in Figure 2, marked by the quandle structure $R:V\to S_{3}=\operatorname{Aut}C_{3}$, realizes $V_{R}$. In the following section, we generalize this by showing that all racks are realized by their full Cayley (di)graphs.

### v4

@@ Abstract.
This article lays the foundations for an analogue of geometric group theory that studies actions on graphs by right quasigroups, including racks and quandles. We study markings of graphs that realize racks, and we introduce (di)graph invariants based on such markings. We show that all right quasigroups are realizable by edgeless graphs and complete (di)graphs. Using Schreier (di)graphs, we also characterize Cayley (di)graphs of right quasigroups $Q$ that realize $Q$. In particular, all racks are realizable by their full Cayley (di)graphs. This solves two problems of Valeriy Bardakov. Finally, we give graph-theoretic characterizations of labeled Cayley digraphs of right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
@@ 1. Introduction
1.1. Motivating discussion
[…]
Racks, quandles, and groups are special classes of nonassociative algebraic structures called right quasigroups, which are used to study column Latin squares [Latin], smooth deformations of Lie group structures [diff], and nonassociative generalizations of Hopf algebras [quantum]. To develop an analogue of geometric group theory for racks and quandles, it is therefore natural to study actions on geometric objects by right quasigroups rather than groups. In this paper, we take those geometric objects to be graphs, particularly labeled Cayley digraphs and other (di)graphs obtained from racks.
1.1.1. Racks as symmetries
In 2020, Bardakov [bardakov] introduced a way to construct right quasigroups from markings of graphs by graph automorphisms. Seeking geometric interpretations of racks, Bardakov posed the following questions about which marked graphs realize racks and how they relate to Cayley graphs of racks.
Problem 1.1.
Under what conditions does a marked graph realize a rack or a quandle?
Problem 1.2.
We answer Bardakov’s questions with the following results. In the following, we refer to directed graphs as digraphs and simple undirected graphs as graphs.
@@ 1.1.1. Racks as symmetries
Let $\Gamma$ be a (di)graph with vertex set $V$, and let $R:V\to\operatorname{Aut}\Gamma$ be a marking (resp. q-marking) of $V$. Then the right quasigroup $V^{\Gamma}_{R}$ realized by $\Gamma$ is a rack (resp. quandle) if and only if $R$ is a magma homomorphism from $V^{\Gamma}_{R}$ to $\operatorname{Conj}(\operatorname{Aut}\Gamma)$.
@@
Theorems 6.1 and 6.2.
@@
All racks are realizable by their full Cayley (di)graphs.
Theorem 5.1 answers Problem 1.1. Since all racks are right quasigroups, Proposition 3.9 answers a generalized version of the first question in Problem 1.2. Corollary 6.4 answers the second question in Problem 1.2, while Theorems 6.1 and 6.2 answer generalized forms of the question. As an application of marked graphs, we also introduce two integer invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ of (di)graphs.
@@ 1.1.2. Characterization of labeled Cayley digraphs
Caucal answered Problem 1.3 for left quasigroups, quasigroups [caucal], semigroups, and various classes of monoids, including groups [caucal2]. Note that the answers to Problem 1.3 for left and right quasigroups are distinct because the right-multiplication maps $R_{v}$ of left quasigroups are not necessarily permutations. Chishwashwa et al. addressed a similar question for vertex-labeled Cayley digraphs of unital, fixed point-free right quasigroups [groupoid1]. In this paper, we answer Problem 1.3 for various classes of racks and right quasigroups.
Theorems 7.12 and 7.14.
Let $\mathcal{Q}$ be the class of all labeled digraphs that are deterministic, source-complete, codeterministic, and target-complete. Then $\mathcal{Q}$ is precisely the class of labeled Cayley digraphs of right quasigroups.
Moreover, there exist graph theoretic-conditions on $\mathcal{Q}$ that restrict it to the subclasses of labeled Cayley digraphs of racks and quandles.
@@ 1.2. Structure of the paper
In Section 5, we answer Problem 1.1. As an application, we introduce two rack-theoretic (di)graph invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ with general results for path graphs and cycle graphs.
In Section 6, we answer Problem 1.2 and its analogues for right quasigroups and digraphs.
In Section 7, we define labeled Cayley digraphs of magmas in the sense of Caucal, and we answer Problem 1.3 for right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
@@ 1.3. Notation
Given a positive integer $n\in\mathbb{Z}^{+}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. Denote the symmetric group of $[n]$ by $S_{n}$ with its elements written in cycle notation, and denote the symmetric group of any other set $X$ by $S_{X}$. We also denote the composition of functions $\varphi:V\to W$ and $\psi:W\to X$ by $\psi\varphi$, and we denote the identity map on a set $V$ by $\operatorname{id}_{V}$. For $n\geq 3$, let $D_{n}$ be the dihedral group of order $2n$. Given a subset $S$ of a group $G$, let $S^{-1}:=\{s^{-1}\mid s\in S\}$.
Acknowledgments
I thank the anonymous referees for their careful reading and helpful comments on previous versions of this paper.
@@ 2.1. Magmas
A magma or groupoid is a pair $(V,R)$, denoted by $V_{R}$, where $V$ is a set and $R$ is a mapping from $V$ to the set of functions from $V$ to $V$. For all $v\in V$, we call the map $R_{v}:=R(v)$ a right-multiplication map or right-translation map.
@@
Remark 2.2.
Although the notation $V_{R}$ is nonstandard, we use it because Bardakov [bardakov] formulated his open problems with it.
@@
If we take $G\coloneq\mathbb{Q}$ to be the set of rational numbers in Example 2.3, then the positive rational numbers $\mathbb{Q}^{+}$ are a submagma of $\mathbb{Q}$ but not a right quasigroup.
@@
Remark 2.6.
Some authors define magmas as sets $V$ equipped with a binary operation $\triangleleft:V\times V\to V$; homomorphisms are then defined in the obvious way. For right quasigroups, racks, and quandles, the operation $\triangleleft$ satisfies additional axioms like the right cancellation property.
This convention is equivalent to our convention via the formula
$v\triangleleft w=R_{w}(v).$
[…]
Evidently, magmas and right quasigroups form categories. In particular, we can consider automorphism groups $\operatorname{Aut}V_{R}$ of magmas.
@@ 2.2.2. Preliminary results
An alternative characterization of racks does the heavy lifting in solving Problem 1.1; see Theorem 5.1.
@@
To prove necessity, suppose that $V_{R}$ is a rack. Then $R(V)\subseteq S_{V}$, and Equation (2.1) states that
@@
To prove sufficiency, suppose that $R$ is a magma homomorphism from $V_{R}$ to $\operatorname{Conj}S_{V}$. Similarly to before,
@@ 3. Graph-theoretic preliminaries
We discuss directed and undirected graphs, marked graphs as constructed by Bardakov [bardakov] and Cayley graphs of magmas as introduced by Caucal [caucal]. We also relate the latter to Schreier graphs of group actions. Since we only discuss labeled digraphs in Section 7, we defer defining them until then.
@@ 3.1. Graphs and digraphs
We recall the graph-theoretic constructions of Bardakov from [bardakov]. Like Bardakov, we assume that all undirected graphs are simple, and we do not allow for digraphs to have multiple edges. However, we do allow for digraphs to have loops.
@@
  • A digraph or directed graph $\Gamma$ is a pair $(V,E)$ where $V$ is a set and $E$ is a subset of $V\times V$. We say that $V$ is the vertex set of $\Gamma$, and we say that $E$ is the (directed) edge set or arc set of $\Gamma$, respectively.
@@ 3.2. Marked graphs
Let $n\in\mathbb{Z}^{+}$ be a positive integer, and let $\Gamma$ be the star graph $K_{1,n}$ of order $n+1$; cf. Table 5.1 in Section 5. Then $\operatorname{Aut}\Gamma\cong S_{n}$ acts on $V:=V(\Gamma)$ by permuting the leaves.
@@ 3.2.1. Right quasigroups are realizable
We answer the first question in Problem 1.2. Recall that a digraph $\Gamma=(V,E)$ is called complete if
@@ 3.4.1. Preliminary results
Our solutions to Problem 1.2 can be stated nicely in terms of Schreier (di)graphs; cf. Theorems 6.16.2.
@@ 4. Motivating examples
Figure 4.1 depicts the full Cayley digraph $\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure 4.1. Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.1.
@@
Let $S=\{1\}$. Figure 4.2 depicts the partial Cayley digraph $\Gamma(V_{R},S)$, the full Cayley digraph $\Gamma(V_{R})$, and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure 4.2. Partial and full Cayley digraphs and full Cayley graph of the quandle from Example 4.2.
@@
Nonisomorphic right quasigroups may share the same Cayley graphs and even the same Cayley digraphs. For example, let $V$ be the set $[3]$. Figure 4.3 depicts the full Cayley digraph of the right quasigroup $V_{R}$ defined by
@@
Figure 4.3. Full Cayley digraphs of the two right quasigroups from Example 4.3 and their shared full Cayley graph.
@@
Figure 4.4 depicts the full Cayley digraph $\Gamma:=\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
@@
Figure $4.$ Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.4.
@@
Let $S=\{1\}$. Figure 4.5 depicts the partial Cayley digraph $\Gamma=\Gamma(V_{R},S)$ and its underlying graph $\Gamma_{\mathrm{und}}(V_{R},S)$.
@@
Figure 4.5. Partial Cayley digraph and underlying Cayley graph of the right quasigroup from Example 4.5.
@@ 5. From marked graphs to racks
In this section, we briefly deduce an answer to Problem 1.1. As an application, we introduce two rack-theoretic invariants of graphs.
5.1. Solution to Problem 1.1
@@
Given a (di)graph $\Gamma$, let $\mu_{\mathrm{rack}}(\Gamma)$ (resp. $\mu_{\mathrm{qnd}}(\Gamma)$) be the number of markings of $\Gamma$ that realize racks (resp. quandles). As an application of Theorem 5.1, we compute these numbers for several graphs and discuss how to compute them in general. This is motivated by Problem 1.1.
@@
We provide an implementation of this exhaustive search algorithm in a GitHub repository [code]. With this implementation, we were able to compute $\mu_{\mathrm{rack}}$ and $\mu_{\mathrm{qnd}}$ for complete graphs $K_{n}$ (equivalently, edgeless graphs), star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$; see Table 5.1. We also have the following general results for path graphs and cycle graphs.
@@
Table 5.1. Computations of $(\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}})$ for complete graphs $K_{n}$, star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$.
@@
When $n=3$, five markings of the cycle graph $C_{3}$ realize quandles. In particular, let $V_{R}$ be the quandle from Example 4.2. The full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)\cong C_{3}$ depicted in Figure 4.2, marked by the quandle structure $R:V\to S_{3}=\operatorname{Aut}C_{3}$, realizes $V_{R}$. In the following section, we generalize this by showing that all racks are realized by their full Cayley (di)graphs.